OpenStudy (anonymous):

to focused on all these essays forgot to study for this 50 question test :( write this expression in quadratic form 3x12+20x6+6

OpenStudy (campbell_st):

well x^12 = (x^6)^2 so its 3(x^6)^2 + 20x^6 + 6 hope it helps

OpenStudy (campbell_st):

you may need a substitution like x^6 = a

OpenStudy (anonymous):

i gets harder :(

OpenStudy (anonymous):

thank you so much !!!